9. Consider the following two demand curves for labor: Qd400- 10*W Qd2 400-20*W a) For each of these, compute the elasticity of demand when the wage is $10 per hour. Answer the following questions for each demand curve If the wage goes up by 10%, what is the percentage change in the quantity demanded. If the wage goes up by 10% what is the change in the number of workers demanded b) d) If the wage goes down by $2 per hour, what is the change in the number of workers demanded
Answer #1

a) Wage elasticity of demand = slope of demand function x Wage / Quantity of labor

ed1 = -10 x 10 / (400 - 10 x 10) = -0.33

ed2 = -20 x 10 / (400 - 10 x 20) = -1

b) For first demand function, the % change in quantity of labor demanded = -0.33 x 10 = -3.3%. For second demand function, the % change in quantity of labor demanded = -1 x 10 = -10%.

c) For first demand function, change in number of workers = 3.3% x (400 - 10*10) = 10 workers will be laid off/fired.

For second demand function, change in number of workers = 10% x (400 - 20*10) = 20 workers will be laid off/fired.

d) For first demand function, change in number of workers =(400 - 8*10) - (400 - 10*10) = 20 more workers are hired

For second demand function, change in number of workers =(400 - 8*20) - (400 - 10*20) = 40 more workers are hired